OpenStudy (anonymous):

*Correct Me! Get A Medal ! [Work Is Shown] Max pays a flat rate of $85 for his cell phone. He is also charged $0.15 for every text that he sends. His cell phone bill can be no more than the amount he has budgeted for it each month. Which of the following is not a possible scenario for the amount Max has budgeted for his cell phone? Max sends 40 texts and has budgeted $96 for his cell phone bill. Max sends 60 texts and has budgeted $97 for his cell phone bill. Max sends 80 texts and has budgeted $98 for his cell phone bill.

OpenStudy (anonymous):

Max sends 100 texts and has budgeted $99 for his cell phone <-------

OpenStudy (anonymous):

0.15(x)-99 0.15(100)-99= -84

OpenStudy (anonymous):

I believe you are correct. If he sends 100 texts, he is charged (100)(.15) = $15 plus his flat rate of $85. So his bill is $100 but he only budgeted $99.

OpenStudy (texaschic101):

(1). 85 + 40(.15) < 96 85 + 6 = 91 (possible) (2) 85 + 60(.15) < 97 85 + 9 = 94 (possible) (3) 85 + 80(.15) < 98 85 + 12 = 97 (possible) (4) 85 + 100(.15) < 99 85 + 15 = 100 (not possible) You are correct.......the 4th one is not possible :)

OpenStudy (anonymous):

Im smart af ! Thank you guy's for correcting me !

OpenStudy (texaschic101):

anytime :)